# MNSC-0001-2024.00

### 15 Mar 2026 v1

@@ 4.3 Analysis of the Robust Optimal Strategy with Permanent Price Impact
A bank’s robust optimal strategy and its worst-case scenario on the aggregate early liquidation size of other banks under various levels of permanent price impact when $\gamma^{*}$ is small We use a four-bank system with $i=1$, $L_{1}=\cdots=L_{4}=6$, $z_{1}=4$, $z_{2}=z_{3}=z_{4}=0.6$, $w_{1}=0.3$, $w_{2}=w_{3}=0.5$, and $w_{4}=0.8$. We set $Q(x)=\exp(-0.2x)$ and $s(x)=\mathop{\arg\min}_{s\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x,s)$, where $g_{i}^{\gamma}$ is defined as in (12).
5 Robust Optimal Strategy with Interbank Exposures
@@
We now take into account debt exposures between banks and investigate how they affect the robust optimal strategy. Below, we present the associated model, which extends the framework introduced in CifuentesEtAl2005 and used in many follow-up studies (ChenEtAl2016, AminiEtAl2016c, WeberWeske2017, CapponiBernard2022). Our extension accounts for (i) the possibility of early liquidation and (ii) each bank’s liabilities to external entities beyond the financial system.
@@
Table 5 describes the initial balance sheet of each bank $i$ in an $m$-bank system operating over two periods. The difference from Table 4.1 lies in that interbank assets $\{L_{ji}\}_{j\neq i}$ are added to the asset side, while total liabilities $L_{i}$ are now composed of liabilities to external entities (denoted by $b_{i}>0$) and interbank liabilities $\{L_{ij}\}_{j\neq i}$, i.e., $L_{i}=b_{i}+\sum_{j\neq i}L_{ij}$. Here, $L_{ij}$ denotes the nominal liabilities of bank $i$ owing to bank $j$, with $L_{ii}=0$, for all $i,j\in[m]$. We again assume that all balance sheet components are assumed to remain unchanged over the two periods ($t=1,2$) and all liabilities are cleared at $t=2$.
[…]
@@
Initial balance sheet of bank $i$ with interbank exposures before liquidation Assets Liabilities and Net Worth Liquid assets $e_{i}$ External liabilities $b_{i}$ Interbank assets $L_{ji},j\neq i$ Interbank liabilities $L_{ij},j\neq i$ Illiquid assets $z_{i}$ with price $P$ Net worth $e_{i}+\sum_{j\neq i}L_{ji}+z_{i}P-L_{i}$ The component $L_{i}$ represents bank $i$’s total liabilities, given by $L_{i}=b_{i}+\sum_{j\neq i}L_{ij}$.
@@
It is also assumed that all liabilities are of equal seniority; if a bank’s asset value at clearing is insufficient to cover total liabilities, its remaining assets are distributed to creditors on a pro rata basis. To illustrate this, we introduce the relative liability matrix ${\bf A}=(A_{ij})$, where $A_{ij}=L_{ij}/L_{i}$ represents the proportion of bank $i$’s liabilities to bank $j$. Then, if bank $j$’s total debt repayment is $\pi_{j}$ for all $j\in[m]$, the total payment received by bank $i$ (i.e., the value of its interbank assets) becomes $\sum_{j=1}^{m}A_{ji}\pi_{j}$ (or equivalently, ${\bf A}_{i}^{\top}\bm{\pi}$). Consequently, given that each bank $i$ sells off $x_{i}$ units of illiquid assets at $t=1$, the market price $p_{2}^{\bf A}$ of illiquid assets at $t=2$ and each bank $i$’s clearing payment $\pi_{j}^{\bf A}$ are defined as the solution to the fixed-point equations, using the inverse demand function $Q$ satisfying Assumption 4.1 over two periods: [^6] In this section, we set aside permanent price impact in order to focus exclusively on the impact of interbank exposures.
@@
$\left\{~\begin{aligned} &p_{2}^{\bf A}=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-x_{j}Q\big({\bf 1}^{\top}{\bf x}\big)-{\bf A}_{j}^{\top}\bm{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)\right),\\ &\pi_{i}^{\bf A}=L_{i}\wedge\left(e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf x}\big)+(z_{i}-x_{i})p_{2}^{\bf A}+{\bf A}_{i}^{\top}\bm{\pi}^{\bf A}\right),~~i=1,\ldots,m.\end{aligned}\right$. (15)
@@
Similar to Section 4, the existence and uniqueness of the above solution is guaranteed by AminiEtAl2016c. Each bank $i$’s liquidation size at $t=2$ is then given by
$\psi_{i}^{\bf A}({\bf x})\coloneqq\frac{\left(L_{i}-e_{i}-x_{i}Q({\bf 1}^{\top}{\bf x})-{\bf A}_{i}^{\top}\bm{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{i}-x_{i}\right).$ (16)
@@
$\phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})\coloneqq e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf x}\big)+\psi_{i}^{\bf A}({\bf x})Q\big({\bf 1}^{\top}\bm{\psi}^{\bf A}({\bf x})\big)+{\bf A}_{i}^{\top}\bm{\pi}^{\bf A}.$ (17)
The subsequent analyses proceed in parallel with Section 4; we (i) relax the liquidation size at clearing, (ii) formulate the robust optimization problem (i.e., the counterpart of (9)), (iii) verify the validity of this formulation, and (iv) derive the robust optimal strategy, which is the optimal solution to this problem. Although the theoretical results here mirror those in Section 4, the proofs are nontrivial extensions and differ significantly due to the presence of $\bm{\pi}^{\bf A}$, which is characterized by a fixed-point equation.
Let us denote $\bar{x}^{\bf A}_{i}\coloneqq\psi_{i}^{\bf A}({\bf 0})$ for each $i\in[m]$, which represents bank $i$’s liquidation size at clearing in the absence of early liquidation, and we assume $\bar{{\bf x}}^{\bf A}>{\bf 0}$ since otherwise liquidation is not required for some banks. We then establish a variant of Lemma 4.2 that gives a robust bound of the liquidation size at $t=2$:
Lemma 5.1
@@
This provides an upper bound on the sum of the middle two terms on the right-hand side of (17), reflecting the worst-case scenario in which banks ultimately liquidate the same volume of illiquid assets as they would without early liquidation. Then, analogous to the construction of the second equation in (15), we can characterize the vector of clearing payments associated with this worst-case scenario, denoted by $\bm{\pi}({\bf c}({\bf x}))$, as the solution to the following fixed-point equation:
$\bm{\pi}={\bf L}\wedge\left({\bf e}+{\bf c}({\bf x})+{\bf A}^{\top}\bm{\pi}\right),$ (19)
where ${\bf c}(\cdot)\coloneqq\big(c_{1}(\cdot),\dots,c_{m}(\cdot)\big)^{\top}$ and $\bm{\pi}({\bf c}({\bf x}))$ is unique for any ${\bf x}\in[{\bf 0},\bar{{\bf x}}^{\bf A}]$ by Theorem 2 of EisenbergNoe2001.
@@
where $\Phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})\coloneqq e_{i}+c_{i}({\bf x})+{\bf A}_{i}^{\top}\bm{\pi}\big({\bf c}({\bf x})\big)$ consists of the initial external liquid assets $e_{i}$, the maximum possible liquidation income $c_{i}({\bf x})$, and the worst-case interbank assets ${\bf A}_{i}^{\top}\bm{\pi}({\bf c}({\bf x}))$. In line with Proposition 4.3, we demonstrate the validity of the above formulation in the following proposition, which addresses (Q2) in Section 3 for the case of interbank exposures:
Proposition 5.2 (Validity of (20))
@@
Theorem 5.3 (Robust Optimal Strategy with Interbank Exposures)
@@
Despite the aforementioned technical complications, the simple characterization of the robust optimal strategy is attributed to our finding that the sensitivity of $\bm{\pi}({\bf c}({\bf x}))$ to $\bf x$} is dominated by that of $c_{i}({\bf x})$ under bank $i$’s worst-case scenario. Consequently, the robust optimal strategy is primarily driven by a solution to the problem $\max_{x_{i}\in[0,\bar{x}^{\bf A}_{i}]}\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}c_{i}({\bf x})$. By applying symmetry and the first-order conditions, its solution is found to be $(\bar{x}^{\bf A}_{i}/2,\bar{{\bf x}}^{\bf A}_{-i}/2)$, leading to the above result.
@@ 6 Conclusion
{APPENDICES}
7 Proofs of Theoretical Results in Section 4
We first introduce several inequalities that will be useful in the proofs of our theoretical results. Specifically, for all $x>0,y\geq 0$ with $x+y<z_{tot}$, let $f(x,y)=xQ(x+y)$ where $Q$ is the inverse demand function. Then, under Assumption 4.1, it is straightforward to check that
@@ 7 Proofs of Theoretical Results in Section 4
Proof 7.1
Proof of Lemma 4.2. Let $\bar{p}\coloneqq Q({\bf 1}^{\top}\bar{{\bf x}})$. By (4) and (5), it is easy to check that
@@
Assume by contradiction that $\bar{p}>p^{\gamma}_{2}$, where $p^{\gamma}_{2}$ is defined as in (4). Then, we first observe that $p_{1}\coloneqq Q({\bf 1}^{\top}{\bf x})\geq Q({\bf 1}^{\top}\bar{{\bf x}})=\bar{p}>p^{\gamma}_{2}$ since $Q(\cdot)$ is strictly decreasing by Assumption 4.1(ii). This implies that for any $\gamma\in[0,1]$,
@@
Then, by (23) and Assumption 4.1(ii), we have
@@
[…]
Proof 7.2
Proof of Proposition 4.3. Fix $x_{i}\leq\bar{x}_{i}$ and let ${\bf u}^{\gamma}\coloneqq\mathop{\arg\min}_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$. By the definition of $\bar{{\bf x}}$ and (21a), we observe that $e_{i}+x_{i}Q(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}_{-i})\leq e_{i}+\bar{x}_{i}Q({\bf 1}^{\top}\bar{{\bf x}})\leq L_{i}$. Thus, we have
@@
Hence, $\psi^{\gamma}_{i}({\bf v}^{\gamma})=z_{i}-x_{i}$ by (5). Since $\bar{x}_{i}\leq z_{i}$ by the definition of $\bar{x}_{i}$ and $\psi^{\gamma}_{j}({\bf v}^{\gamma})\leq\bar{x}_{j}-x_{j}$ for all $j$ by Lemma 4.2, we obtain
@@
[…]
Proof 7.3
Proof of Theorem 4.4. Fix $i\in[m]$. For any $x\in[0,\bar{x}_{i}]$ and $s\in[0,\bar{s}_{i}]$, we define
@@
Step 1. Let $\tilde{x}_{i}^{\gamma}$ be the value satisfying (11) with respect to $x$. We establish its existence and uniqueness given $\gamma\geq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. By Assumption 4.1(iv) and (21c), we observe that $g_{i}^{\gamma}(\cdot,\bar{s}_{i})$ is strictly concave on $0,\bar{x}_{i}]$. Also, it is easy to see that (11) holds if and only if
@@
and thus, by (21d), this value is non-negative if and only if $\gamma\geq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. Similarly, if $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$, by Assumption 4.1 (iv), we obtain
@@
If $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$, Assumption 4.1(ii) and (33) lead to the following relationship:
@@
Furthermore, by Assumption 4.1(ii) and (21d), the following holds for all $(x,s)\in[0,\bar{x}_{i}]\times[0,\bar{s}_{i}]$:
@@
Then, by Lemma 1 of Amir2005, $g_{i}^{\gamma}(x,s)$ is submodular in $0,\bar{x}_{i}]\times[0,\bar{s}_{i}]$, i.e.,
@@
Thus, Topkis’s Monotonicity Theorem (see, e.g., Amir2005, Theorem 1) implies that $\inf{\cal S}(x)$ and $\sup{\cal S}(x)$ are non-decreasing in $x$. This leads to the following relationship:
@@
where $\Delta_{0}\coloneqq\tilde{y}+s(\tilde{y})-(y+s(y))>0$ and $u_{0}\coloneqq\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(y+s(y))$, (41) follows from the mean value theorem, (42) is satisfied since $Q$ is strictly decreasing by Assumption 4.1(ii) and $\tilde{y}+s(\tilde{y})<z^{\gamma}+s(z^{\gamma})\leq(\bar{x}_{i}+\bar{s}_{i})/(2-\gamma)$ by (33), (43) holds because $g_{i}^{\gamma}(y,\cdot)$ is convex and $s(y)$ is its minimizer with $s(y)<\bar{s}_{i}$, and (44) stems from the convexity of $Q$ in Assumption 4.1(iv). On the other hand, if $\bar{s}_{i}, s(y)=s(y~)=$$s(\tilde{y})=\bar{s}_{i}\leq s(z^{\gamma})$ since $s(\cdot)$ is non-decreasing and $y<z^{\gamma}$. This implies that $\gamma\geq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$ and $(1-\gamma)\bar{x}_{i}/(2-\gamma)\leq\tilde{x}_{i}^{\gamma}$ by the definition of $s(\cdot)$ and (33), respectively. Hence, we have $y<\tilde{y}\leq z^{\gamma}=\tilde{x}_{i}^{\gamma}$. Furthermore, by (30) and the strict concavity of $g_{i}^{\gamma}(\cdot,\bar{s}_{i})$, we have
@@
where $u_{1}\coloneqq\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(t+s(t))$, (48) is due to (38) and the mean value theorem, (49) is satisfied since $\tilde{t}+s(\tilde{t})>z^{\gamma}+s(z^{\gamma})=(\bar{x}_{i}+\bar{s}_{i})/(2-\gamma)$, (50) follows by the fact that $t>z^{\gamma}=(1-\gamma)\bar{x}_{i}/(2-\gamma)$ and $Q^{\prime}(\cdot)<0$, (51) holds because $g_{i}^{\gamma}(t,\cdot)$ is convex and $s(t)$ is its minimizer with $s(t)<\bar{s}_{i}$, and (52) stems from Assumption 4.1(iv) and the inequalities $\zeta_{2}<\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(t+s(t))<t+s(t)<\zeta_{1}$. On the other hand, if $s(t)=\bar{s}_{i}$, then for any $x\in(t,\bar{x}_{i}]$, we have $x>z^{\gamma}$. By (31) and (21d), if $\gamma\leq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$,
@@
[…]
Proof 7.4
Proof of Theorem 4.5. Fix $i\in[m]$ and define $g_{i}^{\gamma}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(x+s))$ for each $x\in[0,\bar{x}_{i}]$ and $s\in[0,\bar{s}_{i}]$. Then, (9) can be rewritten as

### 17 Mar 2026 v2

@@ 4.3.  Analysis of the Robust Optimal Strategy with Permanent Price Impact
Figure 3: A bank’s robust optimal strategy and its worst-case scenario on the aggregate early liquidation size of other banks under various levels of permanent price impact when $\gamma^{*}$ is small. We use a four-bank system with $i=1$, $L_{1}=\cdots=L_{4}=6$, $z_{1}=4$, $z_{2}=z_{3}=z_{4}=0.6$, $w_{1}=0.3$, $w_{2}=w_{3}=0.5$, and $w_{4}=0.8$. We set $Q(x)=\exp(-0.2x)$ and $s(x)=\mathop{\arg\min}_{s\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x,s)$, where $g_{i}^{\gamma}$ is defined as in (12).
5.  Robust Optimal Strategy with Interbank Exposures
@@
We now take into account debt exposures between banks and investigate how they affect the robust optimal strategy. Below, we present the associated model, which extends the framework introduced in Cifuentes et al. (2005) and used in many follow-up studies (Chen et al., 2016; Amini et al., 2016; Weber and Weske, 2017; Bernard et al., 2022). Our extension accounts for (i) the possibility of early liquidation and (ii) each bank’s liabilities to external entities beyond the financial system.
@@
Table 2 describes the initial balance sheet of each bank $i$ in an $m$-bank system operating over two periods. The difference from Table 1 lies in that interbank assets $\{L_{ji}\}_{j\neq i}$ are added to the asset side, while total liabilities $L_{i}$ are now composed of liabilities to external entities (denoted by $b_{i}>0$) and interbank liabilities $\{L_{ij}\}_{j\neq i}$, i.e., $L_{i}=b_{i}+\sum_{j\neq i}L_{ij}$. Here, $L_{ij}$ denotes the nominal liabilities of bank $i$ owing to bank $j$, with $L_{ii}=0$, for all $i,j\in[m]$. We again assume that all balance sheet components are assumed to remain unchanged over the two periods ($t=1,2$) and all liabilities are cleared at $t=2$.
@@
Table 2: Initial balance sheet of bank $i$ with interbank exposures before liquidation
Assets Liabilities and Net Worth
Liquid assets $e_{i}$ External liabilities $b_{i}$
Interbank assets $L_{ji},j\neq i$ Interbank liabilities $L_{ij},j\neq i$
Illiquid assets $z_{i}$ with price $P$ Net worth $e_{i}+\sum_{j\neq i}L_{ji}+z_{i}P-L_{i}$
The component $L_{i}$ represents bank $i$’s total liabilities, given by $L_{i}=b_{i}+\sum_{j\neq i}L_{ij}$.
@@
It is also assumed that all liabilities are of equal seniority; if a bank’s asset value at clearing is insufficient to cover total liabilities, its remaining assets are distributed to creditors on a pro rata basis. To illustrate this, we introduce the relative liability matrix ${\bf A}=(A_{ij})$, where $A_{ij}=L_{ij}/L_{i}$ represents the proportion of bank $i$’s liabilities to bank $j$. Then, if bank $j$’s total debt repayment is $\pi_{j}$ for all $j\in[m]$, the total payment received by bank $i$ (i.e., the value of its interbank assets) becomes $\sum_{j=1}^{m}A_{ji}\pi_{j}$ (or equivalently, ${\bf A}_{i}^{\top}\boldsymbol{\pi}$). Consequently, given that each bank $i$ sells off $x_{i}$ units of illiquid assets at $t=1$, the market price $p_{2}^{\bf A}$ of illiquid assets at $t=2$ and each bank $i$’s clearing payment $\pi_{j}^{\bf A}$ are defined as the solution to the fixed-point equations, using the inverse demand function $Q$ satisfying Assumption 1 over two periods: [^6] In this section, we set aside permanent price impact in order to focus exclusively on the impact of interbank exposures.
@@
$\left\{~\begin{aligned} &p_{2}^{\bf A}=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-x_{j}Q\big({\bf 1}^{\top}{\bf x}\big)-{\bf A}_{j}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)\right),\\ &\pi_{i}^{\bf A}=L_{i}\wedge\left(e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf x}\big)+(z_{i}-x_{i})p_{2}^{\bf A}+{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}\right),~~i=1,\ldots,m.\end{aligned}\right$. (15)
@@
Similar to Section 4, the existence and uniqueness of the above solution is guaranteed by Amini et al. (2016). Each bank $i$’s liquidation size at $t=2$ is then given by
$\psi_{i}^{\bf A}({\bf x})\coloneqq\frac{\left(L_{i}-e_{i}-x_{i}Q({\bf 1}^{\top}{\bf x})-{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{i}-x_{i}\right).$ (16)
@@
$\phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})\coloneqq e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf x}\big)+\psi_{i}^{\bf A}({\bf x})Q\big({\bf 1}^{\top}\boldsymbol{\psi}^{\bf A}({\bf x})\big)+{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}.$ (17)
The subsequent analyses proceed in parallel with Section 4; we (i) relax the liquidation size at clearing, (ii) formulate the robust optimization problem (i.e., the counterpart of (9)), (iii) verify the validity of this formulation, and (iv) derive the robust optimal strategy, which is the optimal solution to this problem. Although the theoretical results here mirror those in Section 4, the proofs are nontrivial extensions and differ significantly due to the presence of $\boldsymbol{\pi}^{\bf A}$, which is characterized by a fixed-point equation.
Let us denote $\bar{x}^{\bf A}_{i}\coloneqq\psi_{i}^{\bf A}({\bf 0})$ for each $i\in[m]$, which represents bank $i$’s liquidation size at clearing in the absence of early liquidation, and we assume $\bar{{\bf x}}^{\bf A}>{\bf 0}$ since otherwise liquidation is not required for some banks. We then establish a variant of Lemma 1 that gives a robust bound of the liquidation size at $t=2$:
Lemma 2.
@@
This provides an upper bound on the sum of the middle two terms on the right-hand side of (17), reflecting the worst-case scenario in which banks ultimately liquidate the same volume of illiquid assets as they would without early liquidation. Then, analogous to the construction of the second equation in (15), we can characterize the vector of clearing payments associated with this worst-case scenario, denoted by $\boldsymbol{\pi}({\bf c}({\bf x}))$, as the solution to the following fixed-point equation:
$\boldsymbol{\pi}={\bf L}\wedge\left({\bf e}+{\bf c}({\bf x})+{\bf A}^{\top}\boldsymbol{\pi}\right),$ (19)
where ${\bf c}(\cdot)\coloneqq\big(c_{1}(\cdot),\dots,c_{m}(\cdot)\big)^{\top}$ and $\boldsymbol{\pi}({\bf c}({\bf x}))$ is unique for any ${\bf x}\in[{\bf 0},\bar{{\bf x}}^{\bf A}]$ by Theorem 2 of Eisenberg and Noe (2001).
@@
where $\Phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})\coloneqq e_{i}+c_{i}({\bf x})+{\bf A}_{i}^{\top}\boldsymbol{\pi}\big({\bf c}({\bf x})\big)$ consists of the initial external liquid assets $e_{i}$, the maximum possible liquidation income $c_{i}({\bf x})$, and the worst-case interbank assets ${\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c}({\bf x}))$. In line with Proposition 1, we demonstrate the validity of the above formulation in the following proposition, which addresses (Q2) in Section 3 for the case of interbank exposures:
Proposition 3 (Validity of (20)).
@@
Theorem 3 (Robust Optimal Strategy with Interbank Exposures).
@@
Despite the aforementioned technical complications, the simple characterization of the robust optimal strategy is attributed to our finding that the sensitivity of $\boldsymbol{\pi}({\bf c}({\bf x}))$ to $\bf x$} is dominated by that of $c_{i}({\bf x})$ under bank $i$’s worst-case scenario. Consequently, the robust optimal strategy is primarily driven by a solution to the problem $\max_{x_{i}\in[0,\bar{x}^{\bf A}_{i}]}\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}c_{i}({\bf x})$. By applying symmetry and the first-order conditions, its solution is found to be $(\bar{x}^{\bf A}_{i}/2,\bar{{\bf x}}^{\bf A}_{-i}/2)$, leading to the above result.
@@ 6.  Conclusion
Appendix A Proofs of Theoretical Results in Section 4
We first introduce several inequalities that will be useful in the proofs of our theoretical results. Specifically, for all $x>0,y\geq 0$ with $x+y<z_{tot}$, let $f(x,y)=xQ(x+y)$ where $Q$ is the inverse demand function. Then, under Assumption 1, it is straightforward to check that
@@ Appendix A Proofs of Theoretical Results in Section 4
Proof.
Proof of Lemma 1. Let $\bar{p}\coloneqq Q({\bf 1}^{\top}\bar{{\bf x}})$. By (4) and (5), it is easy to check that
@@
Assume by contradiction that $\bar{p}>p^{\gamma}_{2}$, where $p^{\gamma}_{2}$ is defined as in (4). Then, we first observe that $p_{1}\coloneqq Q({\bf 1}^{\top}{\bf x})\geq Q({\bf 1}^{\top}\bar{{\bf x}})=\bar{p}>p^{\gamma}_{2}$ since $Q(\cdot)$ is strictly decreasing by Assumption 1(ii). This implies that for any $\gamma\in[0,1]$,
@@
Then, by (23) and Assumption 1(ii), we have
@@
which completes the proof. ∎
Proof.
Proof of Proposition 1. Fix $x_{i}\leq\bar{x}_{i}$ and let ${\bf u}^{\gamma}\coloneqq\mathop{\arg\min}_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$. By the definition of $\bar{{\bf x}}$ and (21a), we observe that $e_{i}+x_{i}Q(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}_{-i})\leq e_{i}+\bar{x}_{i}Q({\bf 1}^{\top}\bar{{\bf x}})\leq L_{i}$. Thus, we have
@@
Hence, $\psi^{\gamma}_{i}({\bf v}^{\gamma})=z_{i}-x_{i}$ by (5). Since $\bar{x}_{i}\leq z_{i}$ by the definition of $\bar{x}_{i}$ and $\psi^{\gamma}_{j}({\bf v}^{\gamma})\leq\bar{x}_{j}-x_{j}$ for all $j$ by Lemma 1, we obtain
@@
The second statement is straightforward if $\phi^{\gamma}_{i}(x_{i},{\bf u}^{\gamma})=L_{i}$. Otherwise, it holds by (28). ∎
Proof.
Proof of Theorem 1. Fix $i\in[m]$. For any $x\in[0,\bar{x}_{i}]$ and $s\in[0,\bar{s}_{i}]$, we define
@@
Step 1. Let $\tilde{x}_{i}^{\gamma}$ be the value satisfying (11) with respect to $x$. We establish its existence and uniqueness given $\gamma\geq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. By Assumption 1(iv) and (21c), we observe that $g_{i}^{\gamma}(\cdot,\bar{s}_{i})$ is strictly concave on $0,\bar{x}_{i}]$. Also, it is easy to see that (11) holds if and only if
@@
and thus, by (21d), this value is non-negative if and only if $\gamma\geq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. Similarly, if $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$, by Assumption 1 (iv), we obtain
@@
If $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$, Assumption 1(ii) and (33) lead to the following relationship:
@@
Furthermore, by Assumption 1(ii) and (21d), the following holds for all $(x,s)\in[0,\bar{x}_{i}]\times[0,\bar{s}_{i}]$:
@@
Then, by Lemma 1 of Amir (2005), $g_{i}^{\gamma}(x,s)$ is submodular in $0,\bar{x}_{i}]\times[0,\bar{s}_{i}]$, i.e.,
@@
Thus, Topkis’s Monotonicity Theorem (see, e.g., Amir, 2005, Theorem 1) implies that $\inf{\cal S}(x)$ and $\sup{\cal S}(x)$ are non-decreasing in $x$. This leads to the following relationship:
@@
where $\Delta_{0}\coloneqq\tilde{y}+s(\tilde{y})-(y+s(y))>0$ and $u_{0}\coloneqq\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(y+s(y))$, (41) follows from the mean value theorem, (42) is satisfied since $Q$ is strictly decreasing by Assumption 1(ii) and $\tilde{y}+s(\tilde{y})<z^{\gamma}+s(z^{\gamma})\leq(\bar{x}_{i}+\bar{s}_{i})/(2-\gamma)$ by (33), (43) holds because $g_{i}^{\gamma}(y,\cdot)$ is convex and $s(y)$ is its minimizer with $s(y)<\bar{s}_{i}$, and (44) stems from the convexity of $Q$ in Assumption 1(iv). On the other hand, if $\bar{s}_{i}, s(y)=s(y~)=$$s(\tilde{y})=\bar{s}_{i}\leq s(z^{\gamma})$ since $s(\cdot)$ is non-decreasing and $y<z^{\gamma}$. This implies that $\gamma\geq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$ and $(1-\gamma)\bar{x}_{i}/(2-\gamma)\leq\tilde{x}_{i}^{\gamma}$ by the definition of $s(\cdot)$ and (33), respectively. Hence, we have $y<\tilde{y}\leq z^{\gamma}=\tilde{x}_{i}^{\gamma}$. Furthermore, by (30) and the strict concavity of $g_{i}^{\gamma}(\cdot,\bar{s}_{i})$, we have
@@
where $u_{1}\coloneqq\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(t+s(t))$, (48) is due to (38) and the mean value theorem, (49) is satisfied since $\tilde{t}+s(\tilde{t})>z^{\gamma}+s(z^{\gamma})=(\bar{x}_{i}+\bar{s}_{i})/(2-\gamma)$, (50) follows by the fact that $t>z^{\gamma}=(1-\gamma)\bar{x}_{i}/(2-\gamma)$ and $Q^{\prime}(\cdot)<0$, (51) holds because $g_{i}^{\gamma}(t,\cdot)$ is convex and $s(t)$ is its minimizer with $s(t)<\bar{s}_{i}$, and (52) stems from Assumption 1(iv) and the inequalities $\zeta_{2}<\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(t+s(t))<t+s(t)<\zeta_{1}$. On the other hand, if $s(t)=\bar{s}_{i}$, then for any $x\in(t,\bar{x}_{i}]$, we have $x>z^{\gamma}$. By (31) and (21d), if $\gamma\leq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$,
@@
by (30), the definition of $z^{\gamma}$, and the strict concavity of $g_{i}^{\gamma}(\cdot,\bar{s}_{i})$. This suggests that $)g_{i}^{\gamma}(t,s(t))=g_{i}^{\gamma}(t,\bar{s}_{i})>g_{i}^{\gamma}(\tilde{t},\bar{s}_{i})=g_{i}^{\gamma}(\tilde{t},s(\tilde{t}$)). Consequently, the result follows. ∎
Proof.
Proof of Theorem 2. Fix $i\in[m]$ and define $g_{i}^{\gamma}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(x+s))$ for each $x\in[0,\bar{x}_{i}]$ and $s\in[0,\bar{s}_{i}]$. Then, (9) can be rewritten as